The polynomial of degree 4The polynomial of degree 4, P(x) has a root of multiplicity 2 at x = 4 and roots of multiplicity 1 at x = 0 and x = – 2. It g
Answer #1

solution, Polynomail degree - 4 x=4 has a root of multiplicity is 2 = (x-4) x=0 has a root of multiplicity is = (x-0) x-2 has